OpenStudy (anonymous):

can someone solve this for me to find the ineterest rate. 3,981.04= 3,760.19(1+x/365)^5*365 Not sure what i am doing wrong... I get confused when I get to the exponeent and eliminating it.

OpenStudy (tkhunny):

Just algebra. Why are you getting stuck? 3,981.04= 3,760.19(1+x/365)^5*365 1.0587337 = (1+x/365)^1825 1.000031273696 = 1 + x/365 I'm not seeing a problem. Couple more items to clear up.

OpenStudy (anonymous):

hat did you do with the exponenet?

OpenStudy (anonymous):

exponent* @tkhunny

OpenStudy (anonymous):

I need to solve for x....

OpenStudy (tkhunny):

Well, if you can't do it with those last two steps, you are in the wrong class. You need an exponent button on your calculator. 1.0587337 = (1+x/365)^(5*365) (1.0587337)^(1/5) = (1+x/365)^(365) 1.011480110 = (1+x/365)^(365) 1.011480110^(1/365) = (1+x/365) 1.00003127368 = 1 + x/365 Solve away!

OpenStudy (anonymous):

I'm honestly clue less as to what you did.....

OpenStudy (anonymous):

I'm extremely onfused and I just want to cry....

OpenStudy (tkhunny):

4 = x^2 ==> 4^(1/2) = x 12 = x^3 ==> 12^(1/3) = x 17 = x^5 ==> 17^(1/5) = x 5 = x^7 ==> 5^(1/7) = x You DO need to know your exponent rules. 1.75 = x^52 ==> 1.75^(1/52) = x 1.003 = x^365 ==> 1.003^(1/365) = x These are all the same.